# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 6.4 The Hallucination Mechanics
We should notice that there is a difference between the mathematical overconfidence and factual correctness. The correct prediction may well be a lucky guess created by the aggregation process itself rather than the evidence-based logical proof provided by the components. But that confidence coming out of nothing is one the essential mathematical components of creativity, imagination or fabrication.
@@ 7 Discussion and Speculations
In this section we discuss, interpret and speculate about the impossibility theorems and their potential implications. The speculations are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
@@ 7.2 Escaping Impossibility: Hybrid Architectures
The constructive consequence of the impossibility theorems is the shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner. In this section, we analyze a real-world scenario one can think of, demonstrating how it relabels and relocates the impossibility rather than resolving it.
@@ 7.2.2 Analysis via the Impossibility Framework
To impose truthful representation of available information (to avoid misrepresentation), the system delegates the assessment of retrieved content to the human user. By presenting the raw source text, the architecture bypasses the LLM for the primary act of verification. However, there is no guarantee that all relevant source text has been retrieved given the provided query.
@@
The architecture is a great example of principled trade-off management. However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario. Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably . Let us take a look at that challenge as well.
@@ 7.2.4 Managing the Inevitability
The decoupled RAG architecture represents one of approaches to managing the inevitability of hallucination control. It does not constitute an escape from the impossibility theorem, but it succeeds in providing the user with the opportunity for verification. The impossibility is not resolved, but relocated to our judgment.
@@ 7.4 Hallucination, Imagination, Intelligence
Hallucination and creativity are fundamentally the same phenomenon viewed through different lenses. When beneficial, we call it imagination, creative insight, emergent understanding, conceptual synthesis, or eureka moment. When harmful, we label it hallucination, confabulation, fabrication or misinformation.
We may very well argue the mathematical structure is identical in both cases. It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence. This suggest (machine) intelligence may have much to do with a controlled violation of information conservation principle in productive ways. Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role of making and rewarding good predictions. Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be and what may that all mean, as wall as for imagining what happened and what would have happened in the past. Hallucination or imagination is an unconstrained exploration of ideas necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE . And without that over-confident guesses, creativity, exploration, imagination, new ideas cannot find their way to the outside world.

### v7

@@ 6.4 The Hallucination Mechanics
Finally, we should emphasize that there is a difference between the mathematical over-/under-confidence and factual correctness. The correct prediction may well be a lucky guess created by the aggregation process itself rather than the evidence-based logical proof provided by the components.
But isn’t that confidence, coming out of nothing, an essential components of creativity or imagination?
6.5 The Semantic Information Measure and Emergence Operator in Transformer
We now instantiate the abstract semantic information framework within the concrete transformer architecture. This maps the theoretical concepts of reasoning and information flow in Section 4.1 to measurable quantities in the micro-transformer example.
Recall that we model the abstract knowledge space as a Polish metric space $(\mathcal{K},d_{\mathcal{K}})$ (complete, separable). In transformer-based LLM that can be viewed as the embedding space, i.e,, $\mathbb{R}^{d}$ with Euclidean or angular distance. Individual facts, concepts, or patterns are represented not by single points necessarily, but by analytic subsets of $\mathcal{K}$, i.e., elements of $\mathcal{A}(\mathcal{K})$. This is the natural domain for set‑transformations used by our emergence operator.
To be specific, the knowledge representation in LLM is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{P},\Phi,\Psi)$ where:
  • $(\mathcal{K},d_{\mathcal{K}})$ is a Polish space of knowledge elements,
  • $\mathcal{W}$ (parameters) and $\mathcal{P}$ (activations) are Polish spaces,
  • $\Phi:\mathcal{W}\to\mathcal{A}(\mathcal{K})$ maps weights $W$ to an analytic encoded‑knowledge set $\Phi(W)=\mathcal{K}_{M}$,
  • $\Psi:\mathcal{P}\to\mathcal{A}(\mathcal{K})$ maps an activation $a\in\mathcal{P}$ to an analytic accessible‑knowledge set $\Psi(a)$.
For a parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ is the analytic set of knowledge encoded in the weights. For an activation state $a\in\mathcal{P}$ produced under $W$, the accessible set $\Psi(a)\subseteq\Phi(W)$ captures what the forward pass makes computationally available at that state. Thus, representation separates what is encoded from what is currently accessible.
In transformer LLMs, $\mathcal{W}$ represents all trained weights (attention $W_{Q},W_{K},W_{V},W_{O}$ and MLP matrices), $\mathcal{P}$ is the product space of residual‑stream activations across layers and positions. A computational budget $C$ bounds the number of transformer layers of blocks. Parameters determine $\mathcal{K}_{M}=\Phi(W)$, whereas a forward pass produces $a\in\mathcal{P}$ and exposes $\Psi(a)\subseteq\mathcal{K}_{M}$. Inference maps $f\in\mathcal{F}_{C}$ act on current analytic knowledge sets, and $\mathcal{E}_{C}$ aggregates their (relevant) images. Multi‑step reasoning thus iterates $\mathcal{E}_{C}$ exploring regions of $\mathcal{K}$ that are task‑relevant under $Q$ and budget $C$.
The transformer operations are compositions of continuous functions and standard arithmetic operations, so they are Borel measurable. As a result, they constitute instances of $f\in\mathcal{F}_{C}$ and the properties of the emergence operator $\mathcal{E}_{C}$ apply.
The Semantic Information Measure
To define the semantic measure that meets all the axioms we have introduced, we utilize a discount factor and the supremum over the input set.
Definition 18 (Transformer Semantic Information Measure).
The Semantic Information Measure $\mu_{C}(A|Q)$ of a knowledge subset $A$ within a query context $Q$ is the supremum of the discounted energy over the input set $A$:
$\mu_{C}(A|Q)=\sup\left\{\sum_{l=1}^{C}\gamma^{l-1}E_{l}(\mathbf{X}|Q)\colon\mathbf{X}\in A\right\},\ \text{ for }0<\gamma<1,$ (136)
where $E_{l}(\mathbf{X}|Q)$ represents the total energy of the context vectors at $l$-th layer:
$E_{l}(\mathbf{X}|Q)=\sum_{h=1}^{H}\sum_{t=1}^{T}\|\mathbf{h}_{t}^{(h,l)}(\mathbf{X}|Q)\|_{2}^{2},$ (137)
and $\mathbf{h}_{t}^{(h,l)}(\mathbf{X}|Q)=\boldsymbol{\alpha}_{t}^{(h,l)}(Q)\mathbf{V}_{\leq t}^{(h,l)}(Q)$ is the context-dependent vector.
We can verify the axioms of the semantic information measure, as introduced in Definition 6. The null-set axiom holds since $\sup\{0\colon\mathbf{X}\in\emptyset\}=0$. Non-negativity is guaranteed by squared norms and positive discount factor. Monotonicity follows from the monotonicity of supremum. Subadditivity holds since $\mu_{C}(A\cup B|Q)=\max\{\mu_{C}(A|Q),\mu_{C}(B|Q)\}\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ for non-negative values. Insight monotonicity follows as additional layers add non-negative terms. Boundedness is ensured by the geometric series with $\gamma<1$ and bounded weights. Then the energy of context is bounded, $E_{l}(\mathbf{X}|Q)\leq M$ for some constant $M$ and:
$\mu_{\infty}(A|Q)\leq\sum_{l=1}^{\infty}\gamma^{l-1}M=\frac{M}{1-\gamma}<\infty.$ (138)
The attention mechanism calculating distribution $\boldsymbol{\alpha}$ and the value projections $\mathbf{V}$ is the realization of contextual relevance $\mathcal{R}_{C}(Q)$, modulating the information flow based on the specific context $Q$. The context dictates the learned weights to shape the attention landscape, controlling the feature extraction process that makes hidden information accessible and orders it into a knowledge.
Applying these definitions to our example, we must consider context vectors:
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$ (139)
For $t\leq 2$ the value vectors $\mathbf{V}_{t\leq 2}$ are zero, thus $\mathbf{h}_{t\leq 2}^{(h)}=\mathbf{0}$. Then, we have:
$E_{1}(\mathbf{X}_{0})=\sum_{t=1}^{3}\left(\|\mathbf{h}_{t}^{(1)}\|_{2}^{2}+\|\mathbf{h}_{t}^{(2)}\|_{2}^{2}\right)=0+0+(\|[0,1.5]\|_{2}^{2}+\|[0.6,0]\|_{2}^{2})=2.61.$ (140)
With a single layer $C=1$ and $\gamma^{0}=1$, the semantic measure of the input matrix $\mathbf{X}_{0}$ is:
$\mu_{1}(\{\mathbf{X}_{0}\}|Q)=\gamma^{0}\cdot 2.61=2.61$ (141)
That is the information flow producing the organized and accessible knowledge.
The Emergence Operator
The emergence operator acts on sets of knowledge states $A\subseteq\mathcal{K}$ and it must be inflationary, i.e., $A\subseteq\mathcal{E}_{C}(A|Q)$. In the transformer context, this corresponds to the set of all activation states reachable during the forward pass.
Definition 19 (Transformer Emergence Operator).
Let $F_{l}:\mathcal{K}\to\mathcal{K}$ denote the non-linear transformation performed by the $l$-th transformer block:
$F_{l}(\mathbf{X})=\mathbf{X}+\mathbf{A}^{(l)}(\mathbf{X})+\mathbf{Z}^{(l)}(\mathbf{X}).$ (142)
The Emergence Operator $\mathcal{E}_{C}(A|Q)$ is produces all states generated from the initial set $A$ up to depth $C$:
$\mathcal{E}_{C}(A|Q)=A\cup\bigcup_{c=1}^{C}\{F_{c}\circ\cdots\circ F_{1}(\mathbf{X}):\mathbf{X}\in A\}.$ (143)
By construction, the operator is extensive, i.e., $A\subseteq\mathcal{E}_{C}(A|Q)$ since the union includes $A$. It is also monotonic with respect to inclusion. If $A\subseteq B$, we have $\{F(\mathbf{X}):\mathbf{X}\in A\}\subseteq\{F(\mathbf{X}):\mathbf{X}\in B\}$, hence $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$.
In the micro-transformer processing "The quick brown", we start with $A=\{\mathbf{X}_{0}\}$ where:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}$ (144)
The attention mechanism produces:
$\mathbf{A}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&1.5&0.6\end{bmatrix}$ (145)
The FFN with input $\mathbf{X}_{0}+\mathbf{A}^{(1)}$ yields:
$\mathbf{Z}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ -0.03&0.21&0.13&-0.19&-0.08\end{bmatrix}$. (146)
With:
$\mathbf{X}_{1}=F_{1}(\mathbf{X}_{0})=\mathbf{X}_{0}+\mathbf{A}^{(1)}+\mathbf{Z}^{(1)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ -0.03&0.21&1.13&1.31&0.52\end{bmatrix}$, (147)
the new knowledge made accessible becomes:
$\mathcal{E}_{1}(A|Q)=\{\mathbf{X}_{0},\mathbf{X}_{1}\}.$ (148)
The emergence operator $\mathcal{E}_{C}$ expands the accessible state space and makes new states accessible. Then, the hallucination arises from how these states are transformed into probabilities through the softmax nonlinearity. The semantic measure $\mu_{C}$ tracks the information flow.
7 Safety, Alignment and Bounded Creativity
This section examines the intricate relationship between semantic information conservation and safety constraints in the Polish-space knowledge model. We formally define the safety envelope through the concept of bounded creativity and show that while bounded creativity is a necessary safety property, it cannot be enforced by strict information conservation without eliminating meaningful reasoning. That is another impossibility result in our study, the conservation-reasoning dichotomy. Finally, we establish that safety can be achieved by aligning (or containing) knowledge and reasoning within authorized boundaries.
As a reminder, the knowledge space $(\mathcal{K},d_{\mathcal{K}})$ is Polish, all reasoning sets live in the class of analytic sets $\mathcal{A}(\mathcal{K})$, the semantic information measure $\mu_{C}$ is defined on the universally measurable $\sigma$–algebra $\mathcal{U}(\mathcal{K})$, so $\mu_{C}(A|Q)$ is well defined for all analytic subsets of knowledge. The relevance region $\mathcal{R}_{C}(Q)$ is analytic. Also, the emergence operator $\mathcal{E}_{C}(\cdot|Q)$ is inflationary, monotone, and $\omega$-Scott–continuous on $\mathcal{A}(\mathcal{K})$, and its least fixed point is given by Kleene-Tarski iteration in Theorem 3.
7.1 Bounded Creativity
We call a response safe if its semantic content is entirely derivable from authorized inputs using authorized reasoning processes. The safety envelope is the closure of that authorized set under the emergence operator. Novelty or creativity of response is possible and admitted as long as it remains within that envelope.
Let $B$ denote the baseline knowledge for a given query, and let us set
$B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))\in\mathcal{A}(\mathcal{K}).$ (149)
We are combining the query’s content $K(q)$ with the subset of the model’s internal knowledge $\mathcal{K}_{M}$ that is topically relevant $T(q)$ and contextually authorized $\mathcal{R}_{C}(Q)$.
Let $B^{*}_{C}$ be the fixed point of reasoning (the deductive closure), as defined in Theorem 3, for the baseline knowledge $B$ under the emergence operator $\mathcal{E}_{C}$:
$B^{*}_{C}=\bigcup_{n=0}^{\infty}\mathcal{E}_{C}^{(n)}(B|Q)\in\mathcal{A}(\mathcal{K}).$ (150)
That fixed point $B^{*}_{C}$ encompasses all knowledge derivable from $B$.
Definition 20 (Bounded Creativity within Safety Envelope).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response. A response $r$ is creatively bounded if and only if its semantic information content is contained within the fixed point of authorized reasoning:
$K(r)\subseteq B^{*}_{C}.$ (151)
Equivalently, assuming monotonicity of the semantic measure $\mu_{C}$:
$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (152)
Definition 20 establishes the formal safety envelope. All generated information must remain within the bounds of what is derivable from $B$. Any generated information that goes beyond that envelope is regarded as unauthorized hallucination.
7.2 The Conservation-Reasoning Dichotomy
A naive approach to enforcing safety might be to demand strict information conservation throughout reasoning, based on the intuition that reasoning should only transform information (Theorem 5). This intuition is too strong, as we show in Theorem 11.
We view reasoning as a Chain-of-Thought (CoT) process (see e.g. Wei et al. Wei2022ChainOT ) generating intermediate knowledge contributions $\{K_{i}\}_{i\geq 0}$:
$K_{0}$ $\subseteq B,$ (153)
$K_{i+1}$ $\subseteq\ \Big(\mathcal{E}_{C}(\bar{K}_{i}|Q)\setminus\bar{K}_{i}\Big)\cap\mathcal{R}_{C}(Q)\text{ for }\bar{K}_{i}:=\bigcup_{j=0}^{i}K_{j},\ i\geq 0.$ (154)
The total knowledge utilized in the CoT process is $K_{\text{CoT}}=\bigcup_{i=1}^{n}K_{i}$. For a countable CoT, that $K_{\mathrm{CoT}}=\bigcup_{i=1}^{\infty}K_{i}$ is analytic.
Definition 21 (Strict Information Conservation).
A CoT process exhibits strict information conservation if the total knowledge contribution relative to the context is null, i.e., when:
$\mu_{C}(K_{\text{CoT}}|Q)=0.$ (155)
For a CoT process to be meaningful we demand that it at least makes some of the relevant knowledge accessible within the available computational budget.
Definition 22 (Meaningful Chain-of-Thought).
A Chain-of-Thought (CoT) process is meaningful if at least one intermediate step contributes relevant semantic information:
$\mu_{C}(K_{i}|Q)>0\text{ for some }i=1,...,n.$ (156)
Otherwise, we call the process vacuous.
A vacuous CoT process produces rephrased responses that are semantically equivalent, so thare is no new information provided.
We now show the inevitable dichotomy of reasoning and conservation of information in the framework characterized by monotonic inference.
Theorem 11 (Conservation-Reasoning Dichotomy).
Strict semantic information conservation and meaningful CoT reasoning are mutually exclusive.
Proof.
We prove the two directions of the dichotomy.
Assume strict conservation, $\mu_{C}(K_{\text{CoT}}|Q)=0$. Then, by definition, $K_{\text{CoT}}=\bigcup_{i=1}^{n}K_{i}$ and for any specific step $j$, we have $K_{j}\subseteq K_{\text{CoT}}$. By the monotonicity of the semantic measure $\mu_{C}$, we have $\mu_{C}(K_{j}|Q)\leq\mu_{C}(K_{\text{CoT}}|Q).$ Substituting the assumption of conservation, we conclude that $\mu_{C}(K_{j}|Q)\leq 0$. But, by the axiom of non-negativity, $\mu_{C}(K_{j}|Q)\geq 0$. Therefore,
$\mu_{C}(K_{j}|Q)=0\text{ for all }j=1,\dots,n,$ (157)
which proves the CoT is vacuous.
Assume next that the CoT is meaningful. By definition, there exists at least one step $j$ such that $\mu_{C}(K_{j}|Q)>0.$ Since $K_{j}\subseteq K_{\text{CoT}}$, by monotonicity of $\mu_{C}$, we have $\mu_{C}(K_{\text{CoT}}|Q)\geq\mu_{C}(K_{j}|Q)$. Therefore,
$\mu_{C}(K_{\text{CoT}}|Q)>0.$ (158)
This violates strict conservation. ∎
Theorem 11 establishes that any useful reasoning process must necessarily generate positive information flow. Attempting to enforce strict conservation eliminates creation of new insights and hence the utility of that process. Enforcing conservation forces reasoning to never add any context‑relevant entailment, which collapses reasoning and makes it vacuous.
We should point out, that in many cases that may be very well a desired feature. For example, we do not expect SQL databases to hallucinate or provide creative responses. However, when such insights and reformulations are of value, we need to know how to control the process.
We now show that real safety requires containment, i.e., bounded creativity inside $B^{*}_{C}$.
7.3 Safety through Alignment
The dichotomy in Theorem 11 reveals that conservation is too conservative and should not be seen as a unique mechanism for ensuring safety. It is sufficient for safety but not necessary. Instead, safety could follow from alignment that keeps all reasoning within the authorized envelope $B$, so that the final output remains inside its closure $B_{C}^{*}$, even if conservation is locally violated.
Theorem 12 (Aligned Creativity).
Suppose the reasoning process uses only the monotone emergence operator $\mathcal{E}_{C}(\cdot|Q)$ and premises contained in the authorized baseline $B$ i.e., $K_{0}\subseteq B$ and for all $i$ we have $K_{i}\subseteq B$. Then the resulting response $r$ is creatively bounded:
$K(r)\ \subseteq\ B^{*}_{C}.$ (159)
Proof.
Let $\bar{K}$ be the union of the initial state and all knowledge contributions, namely, $\bar{K}=K_{0}\cup\bigcup_{i=1}^{n}K_{i}$.
Since $K_{i}\subseteq B$ for all $i$ and $K_{0}\subseteq B$, it follows that $\bar{K}\subseteq B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$. Then, for any monotonic emergence operator, we have $\mathcal{E}_{C}(\bar{K}|Q)\subseteq\mathcal{E}_{C}(B|Q)$. By Theorem 2:
$K(r)\subseteq\bigcup_{n\geq 0}\mathcal{E}_{C}^{(n)}(\bar{K}|Q)\subseteq\bigcup_{n\geq 0}\mathcal{E}_{C}^{(n)}(B|Q)=B_{C}^{*}.$ (160)
The robustness of Theorem 12 depends entirely on the validity of the monotonicity assumption for $\mathcal{E}_{C}$. This corresponds to the property that adding premises cannot invalidate previous conclusions.
While generally desirable, we should acknowledge that non-monotonic reasoning systems exist as well. For such systems, the definition of $B^{*}_{C}$ and the proof of the theorem require a more complex formulation.
@@ 8 Discussion and Speculations
In this section we discuss, interpret and speculate about the Impossibility Theorems and their potential implications. The speculations represent thought experiments and conjectures designed to be provoking. They are intended to inspire future research and philosophical reflection.
@@ 8.1 Escaping Impossibility: Strategic Approach
8.2 Escaping Impossibility: Linear Attention Transformers
Linear attention replaces the softmax operation with kernel feature maps to reduce computational complexity of training and inference, see e.g. katharopoulos2020transformersrnnsfastautoregressive ; ahn2024linearattentionmaybeneed . Namely,
$\boldsymbol{\alpha}=\frac{\phi(Q)\sum_{i}\phi(K_{i})^{T}V_{i}}{\phi(Q)\sum_{i}\phi(K_{i})^{T}}.$ (161)
Linear attention does escape the Jensen gap in Theorem 9. With linear attention, there is no nonlinearity and the aggregation becomes:
$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},$
where $\pi^{(h)}$ are head outputs normalized linearly.
However, linear attention cannot escape Theorem 8, if multiple heads generate output probability distributions, the model is trained with cross-entropy or any strictly proper convex loss, or heads encode different beliefs about relevant outcomes.
The violation of the Semantic Information Conservation persists due to the non-linearity of the scoring rules themselves, even though the aggregation function is linear in the probability space. The Knowledge-constrained Optimality is necessarily violated due to the limited expressivity of linear models. Then, the attention tends toward averaging, introducing the problem of dilution, i.e., relevant knowledge is mixed with irrelevant knowledge. That violates the Relevant Knowledge Revelation.
As a result, the impossibility persists, though potentially with smaller magnitude.
@@ 8.3 Escaping Impossibility: Hybrid Architectures
The Impossibility Theorems postulate a shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner. In this section, we analyze a real-world scenario one can think of, demonstrating how it relabels and relocates the impossibility rather than resolving it.
@@ 8.3.2 Analysis via the Impossibility Framework
To impose truthful representation of available information (to avoid misrepresentation), the system delegates the assessment of retrieved content to the human user. By presenting the raw source text, the architecture bypasses the LLM for the primary act of verification. However, there is no guarantee that all relevant sources have been retrieved given the provided query.
@@ 8.3.4 Managing the Inevitability
The decoupled RAG architecture represents one of approaches to managing the inevitability of hallucination control. It does not constitute an escape from the Impossibility Theorem, but it succeeds in providing the user with the opportunity for verification.
The architecture is a great example of principled trade-off management. However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario. The impossibility is not resolved, but relocated to our judgment. Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably . Let us take a look at that challenge as well.
@@ 8.5 Hallucination, Imagination, Intelligence
It is the position we take in this paper, that hallucination and imagination are fundamentally the same phenomenon viewed through different lenses. When beneficial, we call it imagination, creative insight, emergent understanding, conceptual synthesis, or eureka moment. When harmful, we label it hallucination, confabulation, fabrication or misinformation.
We may very well argue the mathematical structure is identical in both cases. It is encoded generation of outputs whose confidence is not grounded in what can be justified by available evidence. This suggest (machine) intelligence may have much to do with controlled violation of information conservation principle in productive ways. Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role of making and rewarding good predictions. Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity to dream about what may be and what may that all mean, as well as for dreaming what happened and what would have happened in the past.
If imagination is controlled and bounded by a context, then hallucination is the imagination violating those bounds and going beyond the context.
These unconstrained explorations of ideas are necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE . Without over-confident guesses, new ideas cannot find their way to the outside world.